Check the convergence of the sequence defined by 𝑒𝑛+1 = (1 + 1/ 𝑒𝑛 ) , 𝑒1 > 0. Note that this is the sequence associated with the continued fraction expansion of the Golden ratio.Β 


1
Expert's answer
2021-12-20T16:45:49-0500

Solution:

Given sequence, "u_{n+1}=1+\dfrac1{u_n},u_1>0"

Let the sequence is convergent to "l".

"\therefore l=1+\dfrac 1l
\\ \Rightarrow l^2=l+1
\\ \Rightarrow l^2-l-1=0
\\ \Rightarrow l=\dfrac{1\pm\sqrt5}{2}
\\ \because u_1>0
\\\therefore l=\dfrac{1+\sqrt5}{2}"

Hence, the given sequence is convergent to "\dfrac{1+\sqrt5}{2}" .